Problem 8-25 Calculating Project NPV

You have been hired as a consultant for Pristine Urban-Tech Zither, Inc. (PUTZ), manufacturers of fine zithers. The market for zithers is growing quickly. The company bought some land three years ago for $1.46 million in anticipation of using it as a toxic waste dump site but has recently hired another company to handle all toxic materials. Based on a recent appraisal, the company believes it could sell the land for $1,560,000 on an aftertax basis. At the end of the project, the land could be sold for $1,660,000 on an aftertax basis. The company also hired a marketing firm to analyze the zither market, at a cost of $131,000. An excerpt of the marketing report is as follows:

The zither industry will have a rapid expansion in the next four years. With the brand name recognition that PUTZ brings to bear, we feel that the company will be able to sell 4,400, 5,300, 5,900, and 4,800 units each year for the next four years, respectively. Again, capitalizing on the name recognition of PUTZ, we feel that a premium price of $710 can be charged for each zither. Since zithers appear to be a fad, we feel at the end of the four-year period, sales should be discontinued.

PUTZ feels that fixed costs for the project will be $455,000 per year, and variable costs are 15 percent of sales. The equipment necessary for production will cost $4.1 million and will be depreciated according to a three-year MACRS schedule. At the end of the project, the equipment can be scrapped for $430,000. Net working capital of $131,000 will be required immediately and will be recaptured at the end of the project. PUTZ has a 40 percent tax rate and the required return on the project is 13 percent. Refer to Table 8.3.

What is the NPV of the project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
NPV           $
Answer #1
Cost of hiring marketing firm is a sunk cost and not relevant for calculation of NPV
Initial Cash Flow:
Opportunity cost of land $1,560,000
Cost of equipment $4,100,000
Net working capital $131,000
Total $5,791,000
A B C=A*B D=C*15% E F=C-D-E G=(1-0.4)*F
Sale in Price per Sales in Variable Fixed cost Before tax After Tax
Year Units unit dollars cost Profit Profit
1                   4,400 $710 $3,124,000 $468,600 $455,000 $2,200,400 $1,320,240
2                   5,300 $710 $3,763,000 $564,450 $455,000 $2,743,550 $1,646,130
3                   5,900 $710 $4,189,000 $628,350 $455,000 $3,105,650 $1,863,390
4                   4,800 $710 $3,408,000 $511,200 $455,000 $2,441,800 $1,465,080
A B=A*$4.1million C=B*40% Tax Rate=40%
Depreciation Amount of Depreciation
Year Rate Depreciation Tax shield
1 33.33% $1,366,530 $546,612
2 44.45% $1,822,450 $728,980
3 14.81% $607,210 $242,884
4 7.41% $303,810 $121,524
After Tax Salvage Value
Before Tax Salvage Value $430,000
After Tax Salvage Value $258,000 (430000*(1-0.4) Present Value of Cash Flow=(Cash Flow)/((1+i)^A)
Terminal Cash Flow: i=Required Return =13%=0.13
Sale of land $1,660,000 A=Year of Cash Flow
Sale of equipment $258,000
Release of working capital $131,000 A B C D E F=B+C+D+E G=E/(1.13^A)
Total terminal cash inflow $2,049,000 initial cash After Tax Depreciation Terminal Net Present value
Year flow Profit Tax shield Cashflow Cash flow PV of cashflow
0 ($5,791,000) ($5,791,000) ($5,791,000)
1 $1,320,240 $546,612 $1,866,852 $1,652,081.42
2 $1,646,130 $728,980 $2,375,110 $1,860,059.52
3 $1,863,390 $242,884 $2,106,274 $1,459,753.54
4 $1,465,080 $121,524 $2,049,000 $3,635,604 $2,229,784.02
TOTAL $1,410,678.49
NPV of the Project $1,410,678.49